# MHONGOOSE — A MeerKAT Nearby Galaxy H i Survey

### v1

@@ 1 Introduction
The evolution of the baryonic matter in the Universe can to a large degree be described as the gradual transformation of primordial atomic hydrogen into galaxies over cosmic time. This transformation involves physical processes on many scales, from the size of galaxy clusters to those of individual gas clouds within a galaxy. These processes, such as gas infall, collapse of clouds, formation of stars, feedback due to stellar winds and supernovae, form part of the “baryon cycle”. Although many of them act on sub-galactic scales, they nevertheless affect the evolution of galaxies as whole.
@@
The only place where a comprehensive detailed survey of these processes can be made is the nearby Universe: locally we can study, in detail, all aspects of the baryon cycle. Resolved observations of the cold gas, and, specifically of neutral hydrogen (H i) can contribute to determining how galaxies acquire their gas, how star formation is sustained, and ultimately how the dark and visible matter interact to determine and regulate the evolution of galaxies.
@@
We concentrate here on the first issue of the origin of gas in galaxies. This question first arose from the observation that in the inner regions of nearby spiral galaxies, time scales for consumption of gas by star formation are much shorter than a Hubble time, even though the star formation rate (SFR) has been approximately constant over most of the lifetime of the galaxies (e.g., Kennicutt 1998; Bigiel et al. 2011; Leroy et al. 2013; Fraternali & Tomassetti 2012). If galaxies are to continue forming stars at their current rate beyond the current epoch, there must be a gas supply external to galaxies. A similar argument can also be made looking at high-redshift galaxies (e.g., Saintonge et al. 2013; Tacconi et al. 2018). These form stars at a rate much higher than observed in the local Universe. This would imply a strong decrease in the gas content of galaxies between then and now, yet we observe an almost constant density, again implying there is an external supply of gas (see, e.g., Walter et al. 2020). In other words, we observe that the gas depletion time for high-redshift galaxies is $\sim 1$ Gyr or less (Saintonge et al. 2013; Tacconi et al. 2018) and yet most galaxies are not quenched and keep forming stars for a much longer time until today.
@@
Galactic disks can, in principle, be replenished by accreting gas-rich companion galaxies, but the slope of the H i mass function is not steep enough for small companions to supply larger galaxies with a substantial amount of gas for a sufficiently long time (Sancisi et al. 2008). In addition, observations suggest that gas-rich mergers can provide at most  20% of the gas required to maintain the SFR in nearby late-type galaxies (Di Teodoro & Fraternali 2014). This implies that spirals have to accrete directly from the intergalactic medium (IGM).
@@
The presence of cold gas in the halos of our Milky Way and other galaxies has been known for several decades (see, e.g., Wakker & Woerden 1997; Oosterloo et al. 2007; Sancisi et al. 2008; Heald et al. 2011 and the reviews by Putman et al. 2012 and Tumlinson et al. 2017). H i in some halos may be part of a star formation-driven “galactic fountain” (Shapiro & Field 1976). This galactic fountain has been proposed to lead to an “indirect” cold gas accretion mechanism, where the gas that is expelled from the disk, drags additional halo gas along as it returns to the disk (Marasco et al. 2012; Fraternali 2017; Marasco et al. 2019).
@@
This is suggested, for example, for the galaxy NGC 2403 by the observation that most of its extra-planar H i has a similar projected radial distribution as the star formation in the disk and that it has disk-like kinematics: rotating but lagging behind the main disk (see, e.g., Fraternali et al. 2001). Similar conclusions were reached for nearby galaxy NGC 253 (Lucero et al. 2015). Note, though, that in NGC 2403, we also observe H i in the halo that likely has an external origin (Veronese et al. 2023), and indeed some of the H i complexes found outside of the main H i disks of galaxies are counter-rotating with respect to the disk, confirming that they cannot have originated from it. Numerical simulations (e.g., Kereš et al. 2005; Ramesh et al. 2023) predict that filaments of cooler gas from the IGM can penetrate the hot halos surrounding galaxies and deposit gas onto the disk. This process is called “cold accretion”. These filaments are a prediction of high-resolution cosmological hydrodynamical simulations of structure formation (e.g., Davé et al. 1999; Crain et al. 2016; Ramesh et al. 2023). These suggest that most of the baryons at low redshift are in a warm-hot ($T\sim 10^{5}-10^{7}$ K) intergalactic medium and most of the gas in the cosmic web is therefore ionised and difficult to observe directly. To detect the smaller fraction of cooler ($T<10^{4}$ K) baryons in the cosmic web, an H i column density sensitivity of $\sim 10^{17-18}$ cm-2 is required (Popping et al. 2009).
It is in the context of this cold accretion that the study of H i halos of galaxies is relevant: it could provide direct observations of the accretion of gas onto galaxies as well as a strong observational test for models of galaxy evolution.
@@
An extensive study of the H i halos of nearby spiral galaxies was made by the Hydrogen Accretion in LOcal GAlaxieS (HALOGAS) project on the Westerbork Synthesis Radio Telescope (WSRT; Heald et al. 2011). For it, 22 disk galaxies were mapped down to an H i column density limit of $\sim 10^{19}$ cm-2, i.e., an order of magnitude lower than the surface densities typically found in star-forming H i disks. The results of HALOGAS indicate that some galaxies have more extended low column density emission, while others do not: extended H i distributions have been detected in about 12 of the 22 galaxies observed. It is possible that some of this gas is related to star formation and galactic fountain processes (Marasco et al. 2019), but external accretion cannot be excluded. The average cold neutral hydrogen gas accretion rate for the HALOGAS galaxies is between 0.05 and 0.09 $M_{\odot}$ yr-1 (Kamphuis et al. 2022), with the exact value depending on the treatment of H i that is detected by the Green Bank Telescope (GBT) but not by the WSRT. This accretion rate is generally lower than the SFR (cf. Table 4 in Kamphuis et al. 2022). If in the HALOGAS sample, the H i accretion rate balances the star formation rate, direct accretion must occur at much lower neutral gas column densities than detected by HALOGAS and on spatial scales not resolved by the GBT.
@@
Very deep single-dish H i observations show the presence of H i at these very low-column densities. Braun & Thilker (2004), for example, observed low-column density features around and between M31 and M33 and, using the WSRT as a single dish, reached a $3\sigma$ limit over 16 km s-1 of $1.1\cdot 10^{17}$ cm-2, but with an angular resolution of $\sim 49^{\prime}$. Further deep observations with the GBT of the HALOGAS (Heald et al. 2011) and The H i Nearby Galaxy Survey (THINGS; Walter et al. 2008) galaxies are described in Pisano (2014), de Blok et al. (2014) and Pingel et al. (2018). These reach a $3\sigma$ over 16 km s-1 sensitivity of $\sim 6\cdot 10^{17}$ cm-2 at an angular resolution of about $9^{\prime}$ or $\sim$ 14 kpc and $\sim 24$ kpc at the median distances of, respectively, the THINGS and HALOGAS galaxies. However, sheer column density sensitivity is not enough, as these relatively coarse spatial resolutions already suggest. Wolfe et al. (2013, 2016) show that the diffuse $\sim 10^{17}$ cm-2 low column density gas between M31 and M33 observed by Braun & Thilker (2004) is resolved in several kpc-sized clouds with peak column densities of a few times $10^{18}$ cm-2 when observed at higher spatial resolutions.
@@
To properly detect and characterise accretion features both a kpc-scale (or better) spatial resolution and a column density sensitivity of $\sim 10^{18}$ cm-2 (or better) are thus needed. Surveys such as THINGS (Walter et al. 2008) and HALOGAS (Heald et al. 2011), have concentrated on either obtaining a high spatial resolution or a high column density sensitivity. Optimising both simultaneously is rarely an option; this has so far limited our knowledge of how any low column density gas is connected with the cosmic web and where accretion occurs.
@@ 2 The MHONGOOSE survey
The array design of MeerKAT is such that the noise level for observations producing beam sizes between $\sim 6^{\prime\prime}$ and $\sim 90^{\prime\prime}$ is approximately flat (see Fig. 1). MeerKAT is therefore not optimised for one particular resolution, but allows high-quality imaging over a large range in resolutions using weighting or tapering.
Thanks to the combination of exquisite column density sensitivity, high spatial resolution (down to $\sim 7^{\prime\prime}$ for H i) and a large field of view with a primary beam full width at half-maximum (FWHM) diameter of 1, we can study nearby galaxies in H i at the required quality to characterise any low-column density H i that may be accreting onto a galaxy.
MHONGOOSE (MeerKAT H i Observations of Nearby Galactic Objects: Observing Southern Emitters) is a MeerKAT Large Survey Project designed to produce ultra-deep H i observations of 30 nearby gas-rich spiral and dwarf galaxies in order to detect and characterise any low-column density, potentially infalling, atomic gas, and to probe its link to star formation (see also de Blok et al. 2016, 2020). These deep observations can provide information on gas flows into and out of the galaxy disks, accretion from the IGM, the fuelling of star formation, the connection with the cosmic web and even the possible existence of low-mass cold dark matter (CDM) halos. The relation between dark and baryonic matter and the distribution of dark matter within galaxies can also be comprehensively studied due to the high resolution combined with high sensitivity. The large field of view of MeerKAT means a significant fraction of the virial volumes of the target galaxies can be observed.
@@
Figure 1: Noise level as a function of resolution for the HALOGAS, THINGS and MHONGOOSE surveys. The blue curve shows that the MHONGOOSE survey has a flat noise distribution, meaning equally sensitive imaging can be obtained over a large range of resolutions between $\sim 6^{\prime\prime}$ and $\sim 90^{\prime\prime}$. The dashed brown curve shows the HALOGAS survey, with an optimal sensitivity around $\sim 20^{\prime\prime}$. The dashed red curve represents the THINGS survey. This is a combination of three separate VLA array configurations (also shown at the top of the plot calculated using the correct relative observing times). THINGS has optimal sensitivity around $\sim 15^{\prime\prime}$ and $\sim 50^{\prime\prime}$. All noise values are calculated over a 5 km s-1 channel.
2.1 Comparison with other surveys
@@
In Fig. 2 we show the sensitivities of these H i surveys and compare these with the observed sensitivity of MHONGOOSE. It is clear that at all resolutions the MHONGOOSE data are significantly more sensitive than previous interferometric observations. Note that the sensitivity does not change simply as the inverse of the beam size squared (which would be a straight line with a slope of $-2$) as one might expect. This is due to the different weightings and taperings used to produce the data at the various resolutions. We will return to this in Sec. 5.6.
@@ 3 Sample Selection
This SINGG/SUNGG proto-sample was divided in bins of 0.2 dex in $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}})$ and in each bin the closest 30 to 40 galaxies were selected. This gave a flat number distribution in the range $10.58.5<\log(M_{{\rm H{\hskip 0.56917pt\tt I}}}/M_{\rm\odot})<$ 10.5, with an additional, small number of galaxies around $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}}/M_{\rm\odot})\sim 7.0$ and $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}}/M_{\rm\odot})\sim 11$. Most of these selected galaxies have a radial velocity $<2000$ km s-1 with a median velocity of 1300 km s-1. These criteria resulted in the final SINGG sample of 468 galaxies as published in Meurer et al. (2006). The key characteristic of this selection is that it was done uniformly in $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}})$ in order to guarantee the broadest possible range in H i masses.
@@
As the goal of MHONGOOSE is to characterise the low-column density H i over a large range in H i masses, we aimed for a flat number distribution in $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}})$ over the available mass range $10.56.0<\log(M_{\rm{\rm H{\hskip 0.56917pt\tt I}}}/M_{\rm\odot})<$ 10.5, analogous to SINGG. To compensate for the smaller number of galaxies, we used six bins each with a width of 0.5 dex, except for the lowest-mass bin which we defined as $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}}/M_{\odot})<8.0$.
The final sample size of 30 galaxies was set by the assigned observing time and the desired column density sentivity. This total number implies a total of five galaxies per mass bin. To optimise the selection for the science objectives of the survey, we required that galaxies be in one of three well-defined inclination categories as follows: i) (close to) face-on, ii) (close to) edge-on and iii) inclination (close to) 60 degrees. Face-on allows the best characterization of the morphology of the interstellar medium (ISM), as well as determination of vertical motions. Edge-on allows an unambiguous characterization of the vertical structure of the ISM and possibly the dark matter distribution out of the disk plane. The intermediate inclination range allows (following a limited amount of modeling) determination of both, and can be used to tie results from edge-on and face-on classes together. It is also an optimal inclination for rotation curve measurements and mass-modeling of the galaxy.
@@
This resulted in the final sample of 30 galaxies. These are listed in Table 1 along with some fundamental properties. Note that some properties listed here are different from those given in the earlier sample table in de Blok et al. (2020). That table lists the sample with the galaxy parameters originally used for the sample selection as based on parameters and distances from SINGG/SUNGG. In the current paper, we have adopted more recent distance estimates for the galaxies (including Tip of the Red Giant Branch [TRGB] measurements). These revised distances resulted in small changes in distance-dependent properties (such as the H i mass), sometimes resulting in a galaxy moving to a different H i mass bin. These changes have no significant impact on the final science goals. The sample still covers the desired large range in H i mass as well as a representative range in (gas-rich) galaxy properties. The current Table lists the galaxies in order of H i mass assuming the revised distances. For the median distance of the galaxies in the MHONGOOSE sample (10.3 Mpc), $10^{\prime\prime}$ corresponds to 0.5 kpc. H i-related parameters listed in the Table are based on the observations presented in this paper. The inclination values are based on the optical axis ratios and should be regarded as indicative only.
@@
In Table 1 we also give the stellar masses and star formation rates based on WISE infrared measurements. The stellar masses were derived using the new GAMA Stellar Masses calibration and (light and colors) method of Jarrett et al. (2023). The star formation rates are a combination mid-IR and UV SFR based on the method described in Cluver et al. (in prep.). Figure 3 shows the distribution of the MHONGOOSE galaxies along the SFR-$M_{\star}$ main sequence (defined by the upper ridges in the diagrams). As the sample was selected to be representative of gas-rich, star-forming galaxies, it therefore covers the main sequence well. The only galaxy not on it is NGC 1371 (J0335–24) which has a low SFR for its stellar mass.
@@
Figure 3: Top panel: Star formation rate of the MHONGOOSE galaxies plotted against their stellar mass. The SFR is derived using the method described in Cluver et al. (in prep.) based on mid-IR and UV SFRs. The stellar masses are derived from WISE W3 and W4 luminosities as described in Jarrett et al. (2023). Background grayscale and gray contours show the distribution of galaxies in the local Universe based on mid-IR and UV data (Jarrett & Cluver, in prep.). The MHONGOOSE sample was selected to be representative of the SFR-$M_{\star}$ main sequence (the upper ridge of the distributions) and almost all galaxies are indeed on this sequence. Only NGC 1371 (J0335–24) has a low SFR for its mass. Galaxies with uncertain SFR values are not plotted (see Tab. 1). UGCA320 (J1303–17b) has a low signal-to-noise in the WISE observations as indicated by the diamond symbol. The purple dashed curve shows the average trend derived in Cluver et al. (in prep.). Bottom panel: Specific SFR of the MHONGOOSE galaxies plotted against stellar mass. Symbols and curves as in top panel. In addition, the blue curve shows a fit to the data presented in Leroy et al. (2019). The diagonal gray dashed lines indicate corresponding SFR values in $M_{\odot}$ yr-1.
@@
Table 1: Basic properties of galaxies in the MHONGOOSE sample
@@
111 Columns: (1) HIPASS identification. (2) Common name. (3) Right ascension (J2000.0). (4) Declination (J2000.0). (5) Distance. Distances taken from Leroy et al. (2019), unless noted in the notes below. (6) Central velocity (this work). (7) Logarithmic H i mass (this work). (8) Indicative inclination (NED). (9) $W_{20}$ H i velocity width (this work). (10) WISE-derived logarithmic stellar mass (11). WISE-derived logarithmic SFR.

### v2

@@ 1 Introduction
The evolution of the baryonic matter in the Universe can to a large degree be described as the gradual transformation of primordial atomic hydrogen into galaxies over cosmic time. This transformation involves physical processes on many scales, from the size of galaxy clusters to those of individual gas clouds within a galaxy. These processes, such as gas infall, collapse of clouds, formation of stars, feedback due to stellar winds, and supernovae, form part of the baryon cycle. Although many of them act on subgalactic scales, they nevertheless affect the evolution of galaxies as a whole.
@@
The only place where a comprehensive detailed survey of these processes can be made is the nearby Universe: locally we can study all aspects of the baryon cycle in detail. Resolved observations of the cold gas, and, specifically of neutral hydrogen (H i) can contribute to determining how galaxies acquire their gas, how star formation is sustained, and ultimately how the dark and visible matter interact to determine and regulate the evolution of galaxies.
@@
We concentrate here on the first issue, that is, the origin of the gas in galaxies. This question first arose from the observation that in the inner regions of nearby spiral galaxies, timescales for the consumption of gas by star formation are much shorter than a Hubble time, even though the star formation rate (SFR) has been approximately constant over most of the lifetime of the galaxies (e.g., Kennicutt 1998; Bigiel et al. 2011; Leroy et al. 2013; Fraternali & Tomassetti 2012). If galaxies are to continue forming stars at their current rate beyond the current epoch, there must be a gas supply external to galaxies. A similar argument can also be made based on observations of high-redshift galaxies (e.g., Saintonge et al. 2013; Tacconi et al. 2018). These form stars at a much higher rate than that observed in the local Universe. This would imply a strong decrease in the gas content of galaxies between then and now; yet we observe an almost constant density, again implying there is an external supply of gas (see, e.g., Walter et al. 2020). In other words, we observe that the gas depletion time for high-redshift galaxies is $\sim 1$ Gyr or less (Saintonge et al. 2013; Tacconi et al. 2018) and yet most galaxies are not quenched and keep forming stars for a much longer time until today.
@@
Galactic disks can, in principle, be replenished by accreting gas-rich companion galaxies, but the slope of the H i mass function is not steep enough for small companions to supply larger galaxies with a substantial amount of gas for a sufficiently long time (Sancisi et al. 2008). In addition, observations suggest that gas-rich mergers can, at most, provide 20% of the gas required to maintain the SFR in nearby late-type galaxies (Di Teodoro & Fraternali 2014). This implies that spirals have to accrete directly from the intergalactic medium (IGM).
@@
The presence of cold gas in the halos of our Milky Way and other galaxies has been known for several decades (see, e.g., Wakker & Woerden 1997; Oosterloo et al. 2007; Sancisi et al. 2008; Heald et al. 2011 and the reviews by Putman et al. 2012 and Tumlinson et al. 2017). The H i in some halos may be part of a star-formation-driven “galactic fountain” (Shapiro & Field 1976). This galactic fountain has been proposed to lead to an “indirect” cold-gas-accretion mechanism, where the gas that is expelled from the disk, drags additional halo gas along as it returns to the disk (Marasco et al. 2012; Fraternali 2017; Marasco et al. 2019).
@@
This is suggested, for example, for the galaxy NGC 2403 by the observation that most of its extraplanar H i has a similar projected radial distribution to the star formation in the disk and that it has disk-like kinematics: rotating but lagging behind the main disk (see, e.g., Fraternali et al. 2001). Similar conclusions were reached for nearby galaxy NGC 253 (Lucero et al. 2015). However, we note that in NGC 2403, we also observe H i in the halo that likely has an external origin (Veronese et al. 2023), and indeed some of the H i complexes found outside of the main H i disks of galaxies are counter-rotating with respect to the disk, confirming that they cannot have originated from it. Numerical simulations (e.g., Kereš et al. 2005; Ramesh et al. 2023) predict that filaments of cooler gas from the IGM can penetrate the hot halos surrounding galaxies and deposit gas onto the disk. This process is called cold accretion. These filaments are a prediction of high-resolution cosmological hydrodynamical simulations of structure formation (e.g., Davé et al. 1999; Crain et al. 2016; Ramesh et al. 2023), which suggest that most of the baryons at low redshift are in a warm-hot ($T\sim 10^{5}-10^{7}$ K) intergalactic medium and most of the gas in the cosmic web is therefore ionized and difficult to observe directly. To detect the smaller fraction of cooler ($T<10^{4}$ K) baryons in the cosmic web, an H i column density sensitivity of $\sim 10^{17-18}$ cm-2 is required (Popping et al. 2009).
@@
It is in the context of this cold accretion that the study of H i halos of galaxies is relevant: it could provide direct observations of the accretion of gas onto galaxies as well as a strong observational test for models of galaxy evolution. An extensive study of the H i halos of nearby spiral galaxies was made by the Hydrogen Accretion in LOcal GAlaxieS (HALOGAS) project on the Westerbork Synthesis Radio Telescope (WSRT; Heald et al. 2011). To this end, 22 disk galaxies were mapped down to an H i column density limit of $\sim 10^{19}$ cm-2, which is an order of magnitude lower than the surface densities typically found in star-forming H i disks. The results of HALOGAS indicate that some galaxies have more extended low-column-density emission, while others do not: extended H i distributions have been detected in about 12 of the 22 galaxies observed. It is possible that some of this gas is related to star formation and galactic fountain processes (Marasco et al. 2019), but external accretion cannot be excluded. The average rate at which cold neutral hydrogen gas is accreted by the HALOGAS galaxies is between 0.05 and 0.09 $M_{\odot}$ yr-1 (Kamphuis et al. 2022), with the exact value depending on the treatment of the amount of H i detected by the Green Bank Telescope (GBT) but not by the WSRT. This accretion rate is generally lower than the SFR (cf. Table 4 in Kamphuis et al. 2022). If in the HALOGAS sample, the H i accretion rate balances the star formation rate, direct accretion must occur at much lower neutral gas column densities than detected by HALOGAS and on spatial scales not resolved by the GBT.
@@
Very deep single-dish H i observations show the presence of H i at these very low-column densities. Braun & Thilker (2004), for example, observed low-column-density features around and between M31 and M33 and, using the WSRT as a single dish, reached a $3\sigma$ limit over 16 km s-1 of $1.1\cdot 10^{17}$ cm-2, but with an angular resolution of $\sim 49^{\prime}$. Further deep observations with the GBT of the HALOGAS (Heald et al. 2011) and The H i Nearby Galaxy Survey (THINGS; Walter et al. 2008) galaxies are described in Pisano (2014), de Blok et al. (2014), and Pingel et al. (2018). These reach a $3\sigma$ over 16 km s-1 sensitivity of $\sim 6\cdot 10^{17}$ cm-2 at an angular resolution of about $9^{\prime}$ or $\sim$ 14 kpc and $\sim 24$ kpc at the median distances of the THINGS and HALOGAS galaxies, respectively. However, sheer column density sensitivity is not enough, as these relatively coarse spatial resolutions already suggest. Wolfe et al. (2013, 2016) show that the diffuse $\sim 10^{17}$ cm-2 low-column-density gas between M31 and M33 observed by Braun & Thilker (2004) is resolved in several kiloparsec-sized clouds with peak column densities of a few times $10^{18}$ cm-2 when observed at higher spatial resolutions.
@@
To properly detect and characterize accretion features both a kpc-scale (or better) spatial resolution and a column density sensitivity of $\sim 10^{18}$ cm-2 (or better) are thus needed. Surveys such as THINGS (Walter et al. 2008) and HALOGAS (Heald et al. 2011) have concentrated on either obtaining a high spatial resolution or a high column-density sensitivity. Optimising both simultaneously is rarely an option; this has so far limited our knowledge of how any low-column-density gas is connected with the cosmic web and where accretion occurs.
@@ 2 The MHONGOOSE survey
The array design of MeerKAT is such that the noise level for observations producing beam sizes between $\sim 6^{\prime\prime}$ and $\sim 90^{\prime\prime}$ is approximately flat (see Fig. 1). MeerKAT is therefore not optimized for one particular resolution, but allows high-quality imaging over a large range in resolutions using weighting or tapering.
Thanks to the combination of exquisite column density sensitivity, high spatial resolution (down to $\sim 7^{\prime\prime}$ for H i) and a large field of view with a primary beam full width at half-maximum (FWHM) diameter of 1, we can study nearby galaxies in H i at the required quality to characterize any low-column density H i that may be accreting onto a galaxy.
MHONGOOSE (MeerKAT H i Observations of Nearby Galactic Objects: Observing Southern Emitters) is a MeerKAT Large Survey Project designed to produce ultra-deep H i observations of 30 nearby gas-rich spiral and dwarf galaxies in order to detect and characterize any low-column density, potentially infalling, atomic gas, and to probe its link to star formation (see also de Blok et al. 2016, 2020). These deep observations can provide information on gas flows into and out of the galaxy disks, accretion from the IGM, the fuelling of star formation, the connection with the cosmic web and even the possible existence of low-mass cold dark matter (CDM) halos. The relation between dark and baryonic matter and the distribution of dark matter within galaxies can also be comprehensively studied due to the high resolution combined with high sensitivity. The large field of view of MeerKAT means a significant fraction of the virial volumes of the target galaxies can be observed.
@@
Figure 1: Noise level as a function of resolution for the HALOGAS, THINGS, and MHONGOOSE surveys. The blue curve shows that the MHONGOOSE survey has a flat noise distribution, meaning equally sensitive imaging can be obtained over a large range of resolutions between $\sim 6^{\prime\prime}$ and $\sim 90^{\prime\prime}$. The dashed brown curve shows the HALOGAS survey, with an optimal sensitivity at around $\sim 20^{\prime\prime}$. The dashed red curve represents the THINGS survey. This is a combination of three separate VLA array configurations (also shown at the top of the plot calculated using the correct relative observing times). THINGS has optimal sensitivity at around $\sim 15^{\prime\prime}$ and $\sim 50^{\prime\prime}$. All noise values are calculated over a 5 km s-1 channel.
@@
In Fig. 2 we show the sensitivities of these H i surveys and compare these with the observed sensitivity of MHONGOOSE. It is clear that at all resolutions the MHONGOOSE data are significantly more sensitive than previous interferometric observations. Note that the sensitivity does not change simply as the inverse of the beam size squared (which would be a straight line with a slope of $-2$) as one might expect. This is due to the different weightings and taperings used to produce the data at the various resolutions. We return to this in Sec. 5.6.
@@ 3 Sample selection
This SINGG/SUNGG protosample was divided in bins of 0.2 dex in $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}})$ and in each bin the closest 30 to 40 galaxies were selected. This gave a flat number distribution in the range $10.58.5<\log(M_{{\rm H{\hskip 0.56917pt\tt I}}}/M_{\rm\odot})<$ 10.5, with an additional, small number of galaxies around $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}}/M_{\rm\odot})\sim 7.0$ and $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}}/M_{\rm\odot})\sim 11$. Most of these selected galaxies have a radial velocity $<2000$ km s-1 with a median velocity of 1300 km s-1. These criteria resulted in the final SINGG sample of 468 galaxies as published in Meurer et al. (2006). The key characteristic of this selection is that it was done uniformly in $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}})$ in order to guarantee the broadest possible range in H i masses.
@@
As the goal of MHONGOOSE is to characterize the low-column density H i over a large range in H i masses, we aimed for a flat number distribution in $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}})$ over the available mass range $10.56.0<\log(M_{\rm{\rm H{\hskip 0.56917pt\tt I}}}/M_{\rm\odot})<$ 10.5, analogous to SINGG. To compensate for the smaller number of galaxies, we used six bins each with a width of 0.5 dex, except for the lowest-mass bin, which we defined as $\log(M_{{\rm H{\hskip 0.56917pt\tt I}}}/M_{\odot})<8.0$.
The final sample size of 30 galaxies was set by the assigned observing time and the desired column density sentivity. This total number implies a total of five galaxies per mass bin. To optimize the selection for the science objectives of the survey, we required that galaxies be in one of three well-defined inclination categories as follows: i) (close to) face-on, ii) (close to) edge-on and iii) inclination (close to) 60 degrees. Face-on allows the best characterization of the morphology of the interstellar medium (ISM), as well as determination of vertical motions. Edge-on allows an unambiguous characterization of the vertical structure of the ISM and possibly the dark matter distribution out of the disk plane. The intermediate inclination range allows (following a limited amount of modeling) determination of both, and can be used to tie results from edge-on and face-on classes together. It is also an optimal inclination for rotation curve measurements and mass-modeling of the galaxy.
@@
This resulted in the final sample of 30 galaxies. These are listed in Table 1 along with some fundamental properties. We note that some properties listed here are different from those given in the earlier sample table in de Blok et al. (2020). That table lists the sample with the galaxy parameters originally used for the sample selection as based on parameters and distances from SINGG/SUNGG. In the current paper, we have adopted more recent distance estimates for the galaxies (including Tip of the Red Giant Branch [TRGB] measurements). These revised distances resulted in small changes in distance-dependent properties (such as the H i mass), sometimes resulting in a galaxy moving to a different H i mass bin. These changes have no significant impact on the final science goals. The sample still covers the desired large range in H i mass as well as a representative range in (gas-rich) galaxy properties. The current Table lists the galaxies in order of H i mass assuming the revised distances. For the median distance of the galaxies in the MHONGOOSE sample (10.3 Mpc), $10^{\prime\prime}$ corresponds to 0.5 kpc. H i-related parameters listed in the Table are based on the observations presented in this paper. The inclination values are based on the optical axis ratios and should be regarded as indicative only.
@@
In Table 1 we also give the stellar masses and star formation rates based on WISE infrared measurements. The stellar masses were derived using the new GAMA Stellar Masses calibration and (light and colors) method of Jarrett et al. (2023). The star formation rates are a combination mid-IR and UV SFR based on the method described in Cluver et al. (in prep.). Figure 3 shows the distribution of the MHONGOOSE galaxies along the SFR-$M_{\star}$ main sequence (defined by the upper ridges in the diagrams). As the sample was selected to be representative of gas-rich, star-forming galaxies, it therefore covers the main sequence well. The only galaxy not on it is NGC 1371 (J0335–24), which has a low SFR for its stellar mass.
@@
Figure 3: Star formation rates of the MHONGOOSE galaxies plotted against their stellar masses. Top panel: the SFR as derived using the method described in Cluver et al. (in prep.) based on mid-IR and UV SFRs. The stellar masses are derived from WISE W3 and W4 luminosities as described in Jarrett et al. (2023). Background grayscale and gray contours show the distribution of galaxies in the local Universe based on mid-IR and UV data (Jarrett & Cluver, in prep.). The MHONGOOSE sample was selected to be representative of the SFR–$M_{\star}$ main sequence (the upper ridge of the distributions) and almost all galaxies are indeed on this sequence. Only NGC 1371 (J0335–24) has a low SFR for its mass. Galaxies with uncertain SFR values are not plotted (see Tab. 1). UGCA320 (J1303–17b) has a low S/N in the WISE observations as indicated by the diamond symbol. The purple dashed curve shows the average trend derived in Cluver et al. (in prep.). Bottom panel: this shows the specific SFR of the MHONGOOSE galaxies plotted against stellar mass. Symbols and curves as in top panel. In addition, the blue curve shows a fit to the data presented in Leroy et al. (2019). The diagonal gray dashed lines indicate corresponding SFR values in $M_{\odot}$ yr-1.
@@
Table 1: Properties of galaxies in the MHONGOOSE sample
@@
111 Columns: (1) HIPASS identification. (2) Common name. (3) Right ascension (J2000.0). (4) Declination (J2000.0). (5) Distance. Distances taken from Leroy et al. (2019), unless specified otherwise in the notes below. (6) Central velocity (this work). (7) Logarithmic H i mass (this work). (8) Indicative inclination (NED). (9) $W_{20}$ H i velocity width (this work). (10) WISE-derived logarithmic stellar mass (11). WISE-derived logarithmic SFR.